How many litres of 50% alcohal solution and 20% alcohal solution must be mixed to obtain 18 Liters of 30% alcohal solution?

Answer by stanbon(75887)

How many litres of 50% alcohol solution and 20% alcohol solution must be mixed to obtain 18 Liters of 30% alcohol solution?
----
Equation:
active + active = active
---
0.50x + 0.20(18-x) = 0.30*18
Multiply thru by 100 to get:
50x + 20*18 - 20x = 30*18
30x = 10*18
x = 6 liters (amt. of 50% solution needed
18-x = 12 liters (amt. of 20% solution needed.
=====================================
------